OpenStudy (anonymous):

In the expression (x + y)8, if x = 0.3 and y = 0.7, what is the numerical value of the third term?

OpenStudy (anonymous):

0.010 0.020 0.058 0.060 0.068

OpenStudy (anonymous):

x=0.3 and y=0.7, then 8(x+y) = 8(0.3 +0.7) = 8(1)=8

OpenStudy (anonymous):

there is a feeling you haven't given all the information in the question.

OpenStudy (anonymous):

that is all the information

OpenStudy (anonymous):

your answer dont go with any answer choices

OpenStudy (anonymous):

i think you gotta use the binomial theorem

OpenStudy (anonymous):

Well then you didnt give me all the information, as I suspected

OpenStudy (anonymous):

thats all they gave me to answer it thats why i need help with this

OpenStudy (anonymous):

if you are talking about a third term, then you must be talking about some kind of progression or sequence. You have not provided a sequence(either by closed form or recurrence sequence or difference equation), therefore we cannot calculate the third term.

OpenStudy (anonymous):

i get it you just dont know how to answer it its ok

OpenStudy (anonymous):

Sometimes, the question is bad, and maybe in this case, that is why.

OpenStudy (anonymous):

ok.. ig